OpenStudy (anonymous):

Can someone help me in deriving the equation for the electric field due to a charged disk?

OpenStudy (anonymous):



OpenStudy (vincent-lyon.fr):

Hi guys! Try to work on that by yourselves. I can help later or comment on your formulae.

OpenStudy (anonymous):

equation of electic feild due to a charged disk.

OpenStudy (anonymous):

okay..i did it like this - i just wanna know where i went wrong

OpenStudy (anonymous):

then you have to see it in your work nah..

OpenStudy (anonymous):

sigma( area charge density) = dq/dA or sigma * dA = dq; =sigma * (2 pie r dr) dE = 1/4 pie (ep) * (sigma* 2 pie r dr)/(r^2 + z^2)^2 along x axis, E will be zero along Y axis; dEsin theta comes out to be 1/4pie (ep) * (sigma*2 pie r^2 dr)/(r^2 + z^2)^3/2; (put the value of sin theta as r/ sqrt (r^2 + z^2) and check that; from here, when i integrate i get 1/2 (ep) * sigma r^2 * 2 pie r/(r^2+z^2)^3/2;

OpenStudy (anonymous):

ep = epsilon

OpenStudy (aravindg):

@goutham1995 pls use eqn editor

OpenStudy (anonymous):

@AravindG - how?

OpenStudy (aravindg):

on bottoof reply u have equation draw attach file, click equation then type in

OpenStudy (aravindg):

*bottom

OpenStudy (experimentx):

\[ \int_{0}^{a}\frac{kq\sigma2\pi r dr \cos\theta}{r^2 + x^2}\] \( \cos\theta = \frac{x}{\sqrt{x^2+r^2}}\)

OpenStudy (anonymous):

\[\sigma/ 2\epsilon (1 - z/\sqrt{z^2 + R^2) is the answer

OpenStudy (anonymous):

what--sorry it didnt print out properly -

OpenStudy (anonymous):

sigma/2 (ep)( 1 - z/sqrt(z^2 + R^2)) should be the answer

OpenStudy (anonymous):

\[\sigma/2\epsilon (1 - z/\sqrt{z ^{2 }+R ^{2}}\]